The weekly output (in units) of a factory depends on the amount of capital and labour it employs as follows: if it uses k units of capital (K) and l unit of labour (L) then its output is Q=L^0.5K^0.5 units. The cost to the firm of each unit of capital is $4, and the cost of each unit of labour is $1. Use the method of Lagrange multipliers to find the minimum weekly cost of producing a quantity of 200 units.
1
Expert's answer
2021-10-24T20:16:59-0400

Given that: k units of capital and l units of labour


Q(k, l)="\sqrt{k}\sqrt{l}"


The cost to the firm of each unit of capital is $4, and the cost of each unit of labour is $1.


Q(k, l)="\sqrt{4}\sqrt{1}"


Q(k, l)=(2)(1)=2


"\sqrt{k}\sqrt{l}" =2


P(k, l)="\sqrt{k}\sqrt{l}" -2...........................(say)


"\nabla"Q(k, l)="\frac{1}{2\sqrt{k}},\frac{1}{2\sqrt{l}}"


"\nabla"P(k, l)="\frac{1}{2\sqrt{k}},\frac{1}{2\sqrt{l}}"


By Lagrange multipliers;


"\nabla"Q="\lambda\nabla"P


("\frac{1}{2\sqrt{k}},\frac{1}{2\sqrt{l}}") ="\lambda" ("\frac{1}{2\sqrt{k}},\frac{1}{2\sqrt{l}}")


"\lambda" =1, k=4 and l=1


Q(k, l)="\sqrt{k}\sqrt{l}" =2 (The minimum weekly cost of producing a quantity of 2 units)


The minimum weekly cost of producing 200 units will therefore be:

(200 units"\div" 2 units)"\times 2\$" =200"\$" =200 dollars.